OpenStudy (anonymous):

Which table best represents the steps to solve the equation 8 + x over 3 = 4?

OpenStudy (anonymous):

Step Justification Step 1: x over 3 = 4 + 8 The same number +8 can be added to both sides of the equation Step 2: x over 3 = 12 Adding 8 and 4 on the right hand side Step 3: x = 12 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = 36 Multiplying 12 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = -12 Multiplying -4 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number - 8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding 4 and -8 on the right hand side tep 3: x = - 4 - 3 The same number 3 can be subtracted from both sides of the equation Step 4: x = -7 Simplifying -4 -3 Step Justification Step 1: x over 3 = -4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 ÷ 3 The same number 3 can be divided on both sides of the equation Step 4: x = -1.33 Dividing -4 by 3 on the right hand side

OpenStudy (shinalcantara):

second